# Galaxy clustering in a new implementation of the GAEA semi-analytical galaxy formation model

### v1

@@
Galaxy clustering in a new implementation of the GAEA semi-analytical galaxy formation model
We present results from the latest version of the GAlaxy Evolution and Assembly (gaea ) theoretical model of galaxy formation coupled with merger trees extracted from the P-Millennium Simulation (PMS ). With respect to the Millennium Simulation, typically adopted in our previous work, the PMS provides a better mass resolution ( 10), a larger volume (8003 Mpc3) and assumes cosmological parameters consistent with latest results from the Planck mission. The model includes, at the same time, a treatment for the partition of cold gas into atomic and molecular (H2) components; a better treatment for environmental processes acting on satellite galaxies; an updated modelling of cold gas accretion on Super-Massive Black Hole, leading to the Active Galactic Nuclei (AGN) phenomenon and relative feedback on the host galaxy. We compare gaea predictions based on the PMS , with model realizations based on other simulations in the Millennium Suite at different resolution, showing that the new model provides a remarkable consistency in most statistical properties of galaxy populations. We interpret this as due to the interplay between AGN feedback and H2-based SFR (both acting as regulators of the cold gas content in model galaxies), as model version considering only one of the two mechanisms do not show the same level of consistency in the predictions. We then compare model predictions with available data for the galaxy 2-point correlation function (2pCF) in the redshift range 0¡z3. We show that gaea runs are able to correctly recover the main dependencies of the 2pCF as a function of stellar mass (M⋆), star formation activity, HI-content and redshift for M⋆ ¡ 1011 M⊙ galaxies. These results suggest that our model correctly captures both the distribution of galaxy populations in the Large Scale Structure and the interplay between the main physical processes regulating their baryonic content, both for central and satellite galaxies. At larger stellar masses gaea underpredicts the 2pCF amplitude, suggesting that model massive galaxies live in less massive dark matter haloes. The model predicts a rather small redshift evolution of the clustering amplitude up to z3, consistent with available observational evidence.
@@
According to the classical interpretation of the 2pCF shape, it can be seen as a combination of (see, e.g., Cooray & Sheth, 2002): a 2-halo term - corresponding to pairs of galaxies belonging to independent halos - and a 1-halo term - coming from pairs of galaxies living in the same halo. The 2-halo term is dominant at large scales (separations larger than 3 Mpc) and comes mainly from pairs of central galaxies. The 1-halo term is responsible for the 2pCF shape at 2 Mpc and it is statistically dominated by galaxy pairs belonging to the same halo (that is at least one of the two galaxies is a satellite). Overall, the largest contribution to the global 2pCF comes from galaxies around the knee of the GSMF, i.e. with stellar masses similar to the Milky-Way (Li & White, 2009). The small scale clustering is particularly sensitive to the number and distribution of satellite galaxies inside massive DMHs. Therefore, it provides important constraints to theoretical models of galaxy formation, as it critically depends on the treatment adopted for satellite galaxies and in particular their survival times. Therefore, it can be significantly affected by numerical resolution effects (Guo & White, 2014).
Observationally, only the projected 2pCF - - is directly accessible, while the tridimensional 2pCF can be recovered via de-projection. Early results by Davis & Geller (1976) already showed a clear dependence of clustering amplitude with galaxy type, by showing that early type galaxies are more clustered than late-type galaxies. The advent of large area redshift surveys like the Sloan Digital Sky Survey (SDSS) provided the statistics to explore the dependence of amplitude and shape on galaxy properties such as luminosity (see, e.g., Zehavi et al., 2002) or stellar mass (Li et al., 2006), showing that brighter/more massive galaxies are more clustered than their fainter/less massive counterparts. A stronger difference in the clustering signal is found when considering galaxies of different colour (see, e.g., Li et al., 2006); at fixed M⋆ (or luminosity), red galaxies sensibly more clustered than blue galaxies. Galaxy colour is sensitive to the star formation rate (SFR) of model galaxies, which is regulated by a complex network of internal and external physical mechanisms. Therefore, clustering predictions for the active and passive galaxy populations represent a additional diagnostic to assess whether a given model reproduces the observational data in terms of star formation quenching as a function of the environment.
@@
Most of our previous gaea realizations have been run on dark matter merger trees extracted from the Millennium Simulation (Springel et al., 2005, MS hereafter). This realization is based on a cosmological model whose parameters are offset with respect to more recent measurements – particularly for the normalisation of the density fluctuations at the present day () and . In this paper, we will introduce a novel gaea realization, run on the p-millennium simulation (Baugh et al., 2019, PMS hereafter), featuring a bigger cosmological volume, an updated cosmology, a better mass resolution and a finer time sampling. In particular, we will discuss the redshift evolution of key statistical quantities like the GSMFs or the AGN LFs as well as of galaxy clustering, as a function of stellar mass and star formation activity.
@@ 2 N-body simulations and the galaxy formation model
Star Formation efficiency | ||
@@
rpstime | 400 | Timescale of hot gas ram-pressure stripping |
kesislow | 1.4 | Ratio between specific angular momentum of gas cooling through ‘slow mode’ and that of the halo |
kesirapid | 3.0 | Ratio between specific angular momentum of gas cooling through ‘rapid mode’ and that of the halo |
@@ 2.1 Numerical Simulations
In this work, we will focus on predictions from a gaea realization coupled with merger trees extracted from the p-millennium N-body simulation Baugh et al. (PMS 2019). Moreover, we will compare our results against predictions from two other gaea runs, based on the millennium Simulation (Springel et al., 2005, MS) and millennium ii Simulation (Boylan-Kolchin et al., 2009, MSII). These runs corresponds to the predictions at the basis of our recent work (De Lucia et al., 2024; Xie et al., 2024). Details on the assumed cosmologies, box size and particle resolution for each N-body simulation are listed in Table 1.
@@
The initial conditions for the PMS have been generated at z=127 using second order Lagrangian perturbation theory (see Jenkins, 2013, for more details), and the simulation has been run using a reduced memory version of the N-body code gadget (Springel et al., 2005). Halo and Sub-Halo catalogues have been extracted on the fly using a classical FOF algorithm and the substructure finder code subfind (Springel et al., 2001): halos and sub-haloes information have been saved on a regular grid of 271 output times (i.e. snapshots), equally spaced in expansion factor. This has to be compared with the 64-68 snapshots of the MS and MSII over the same redshift range. The gain in temporal (redshift) resolution is particularly relevant at z4 where the PMS has 100 snapshots against the of the MS/MSII. In order to run GAEA on the PMS volume, we have constructed subhalo based merger trees using the same approach described in Springel et al. (2005). We refer to this paper for a detailed description of the algorithm. In summary, for any given subhalo, we identify a unique descendant by finding all subhalos at the subsequent output that contain its most bound particles. We then give higher weight to those particles that are more tightly bound to the subhalo under consideration, and select as descendant the subhalo containing a larger fraction of the most bound particles. In this way, we aim at tracing the fate of the inner regions of the substructure, that might survive for longer time after it has been accreted onto a larger structure. Following (Springel et al., 2005), we search for descendants in the two following outputs to deal with small subhalos that fluctuate below and above the assumed detection threshold (20 bound particles). Once a unique descendant has been identified for all subhalos in the simulated volume and for all available outputs, links to all progenitors are automatically established, and merger trees can be built. We adopt the same merger tree structure defined for the Millennium Simulation suite and shown in Fig. 11 of (Springel et al., 2005). The merger trees adopted in this work have been constructed using half of the available outputs. While reducing the number of redshifts available, this choice reduces the memory and computational requirements of our GAEA runs. In a forthcoming paper (Cantarella et al., in preparation), we will show that results are not significantly different from those obtained using merger trees constructed considering all available snapshots.
@@ 4.1 Redshift evolution of the 2-point correlation function
We then consider the predicted correlation functions for our model universes at different redshifts. In the following, we will focus on the projected 2-point correlation function ; this choice makes the comparison with observational data straightforward, and does not have to rely on de-projection algorithms. We study the evolution of the projected correlation function using estimates from several surveys: going from low-z to higher redshifts we compare against results from SDSS (Li et al., 2006, z0.1), GAMA (Farrow et al., 2015, z¡0.2), VIPERS (Marulli et al., 2013, ), DEEP2 (Li et al., 2006, z0.9), VVDS (Meneux et al., 2008, z0.9), zCOSMOS (zCOSMOS Meneux et al., 2009, z0.9) and VUDS (Durkalec et al., 2018, z2.7). We collect all these observational measurements in Fig. 3 and we contrast them with the prediction from gaea realizations based on the MS (black dashed lines) or the PMS (red solid lines). We compute from gaea using the python corrfunc package (Sinha & Garrison, 2020); model galaxy samples have been extracted from the simulated volumes considering all galaxies in the stellar mass range indicated in the legend, i.e. we do not attempt to replicate selection criteria specific of the different surveys considered. In our analysis we include an estimate for the typical error on M⋆ by convolving the intrinsic predictions of the model with a log-normal error distribution of amplitude 0.25 dex. We check that our main conclusions do not depend on this choice: as predictions based on intrinsic M⋆ provide a consistent picture.
Overall, the level of agreement between theoretical and observed correlation functions is impressive over the entire redshift range considered. Models run on the MS and PMS agree well with each other, but for a tendency of the MS runs to predict systematically stronger correlations. The effect is more pronounced at small stellar masses and close separations, and we ascribe it to the larger value of used in the MS. At z (Fig. 3, upper left panel), theoretical are very close to SDSS constraints for , but they tend to underestimate the correlation signal measured for more massive galaxies. Similar conclusions hold at intermediate redshifts () for VIPERS (Fig. 3, upper right panel), but the underprediction in the highest mass bin reduces considerably as the redshift increases. The situation is slightly more complex at z0.9, where different surveys suggest slightly different normalization for . This could be due to the different surveyed volume or to the presence of overdensities (like in the case of the zCOSMOS sample, (Meneux et al., 2009)). We explore the impact of cosmic variance by considering 125 independent and disjointed subvolumes in the PMS , and computing their corresponding 2pCF. The yellow areas in Fig. 3 highlight the range covered by these independent realizations: smaller amplitudes are indeed possible, especially at the largest separations. Nonetheless, gaea predictions lie within the clustering measurements obtained from different surveys at z0.9 (Fig. 3, lower left panel) and the agreement with data is good to the smallest separations measured. The predicted evolution of the normalization at z1 is rather small: in most cases can be well approximated by a single power law over a wide separation scale; gaea also reproduces the observed flattening of the relation at low-z for the smallest separations probed by the SDSS. Moving to higher redshifts (z), a larger difference in the ) normalization can be appreciated with respect to the VUDS data (Fig. 3, lower right panel): uncertainties in the observational determination are typically larger than at lower redshift, and gaea predictions are still consistent, although systematically larger, than the data. Cosmic variance can also have a large impact here (the VUDS volume correspond to 1.75 107 Mpc3, to be compared with the 4 106 Mpc3 of the PMS subvolumes). We thus conclude that gaea is are able to correctly capture the observed evolution of the correlation strength up to z3.5: this represents a major success for our latest gaea model release.
@@ 4.2 Colour/SFR dependence of the 2-point correlation function
As we already discussed, several authors (see e.g. Wang et al. 2013) pointed out that the clustering strength depends not only on , but also on the star formation activity of galaxy population, with red and quiescent galaxies being more clustered than their blue and star forming counterparts with similar . In Fig. 4, we compare model predictions with SDSS data at z0.1 (Li et al., 2006) and data at 0.9 (Coil et al., 2017) obtained combining the PRIsm MUlti-object Survey (PRIMUS Coil et al., 2011) with DEEP2. At both redshifts we consider model predictions corresponding to quiescent and star-forming subsamples, using separation thresholds based on their specific SFR (sSFR = SFR/M⋆), for consistency with our previous work (De Lucia et al., 2024). At z0.1, model galaxies are classified using a sSFR= 10-11 yrs-1 separation. This is formally different than the choice for SDSS galaxies, which have been split according to their g-r colour: however, we checked that using a classifications based on galaxy colours (e.g. as in Springel et al. 2018) does not qualitatively change our main results: gaea realization reproduces well the stronger clustering of red galaxies with respect to blue sources, and in particular the relative clustering amplitudes between the two populations. Models over-predict the clustering strength for red galaxies at intermediate masses, irrespectively of the assumed cosmology. Overall, the clustering strength of blue galaxies is better reproduced over the entire mass range considered. At large masses, the model under-predict the observed clustering signal, but the relative clustering between the red and blue population is still well represented. At higher redshift, we contrast clustering estimates from the PRIMUS and DEEP2 surveys with gaea predictions by splitting model galaxies into a quiescent and star forming subsamples, using the same (redshift-dependent) sSFR threshold employed in the data (Coil et al., 2017):
The agreement between data and model prediction in the single mass bin available (Log(M]) ) is very good, both for the red and blue populations. Overall, the results presented in Fig. 4 indicate that the treatment for physical mechanism responsible for quenching in gaea correctly reproduces not only the overall assembly and evolution of the quenched galaxy population (De Lucia et al., 2024; Xie et al., 2024), but also their spatial distribution in the cosmic web. This confirms that the model correctly capture the relative role of internal and environmental processes in shaping the galaxy properties in different environments.
@@
In the previous sections, we study the 2pCF of galaxies selected on the basis of their stellar content. A complementary analysis involves the study of the cold gas content of galaxies, through the distribution of HI selected galaxies. Gas distribution in galaxies at low redshift can be mapped using the 21 cm HI hyper-fine emission line. In particular, the Arecibo Fast Legacy ALFA Survey (ALFALFA; Giovanelli et al. 2005) samples more than 5000 galaxies out to z0.05, enabling the study of their clustering properties Papastergis et al. (2013), showed that HI selected galaxies exhibit a lower amplitude with respect to stellar mass selected samples. Guo et al. (2017) used the 70% complete ALFALFA catalogue and showed that the 2pCF of of HI-selected galaxies depends strongly on the HI mass, i.e. galaxies with higher HI reservoirs are more strongly clustered on scales above a few Mpc than their counterparts with lower HI content. We compare Guo et al. (2017) estimates at different HI-mass (MHI) thresholds with gaea predictions in Fig. 5; we find a general good agreement in the clustering amplitude and, most importantly, in its relative evolution as a function of MHI. The larger discrepancies are seen at the high-MHI end, although data and predictions are consistent within the statistical errorbars. This result is similar to those based on M⋆ selected galaxies, although the reference samples refer to different model galaxy populations, as most massive galaxies at in gaea have low gas fractions.

### v2

# Galaxy Assembly and Evolution in the P-Millennium simulation: galaxy clustering

@@
Galaxy Assembly and Evolution in the P-Millennium simulation: galaxy clustering
We present results from the latest version of the GAlaxy Evolution and Assembly (gaea ) theoretical model of galaxy formation coupled with merger trees extracted from the Planck Millennium Simulation (PMS ). With respect to the Millennium Simulation, typically adopted in our previous work, the PMS provides a better mass resolution ( 10), a larger volume (8003 Mpc3) and assumes cosmological parameters consistent with latest results from the Planck mission. The model includes, at the same time, a treatment for the partition of cold gas into atomic and molecular (H2) components; a better treatment for environmental processes acting on satellite galaxies; an updated modelling of cold gas accretion on Super-Massive Black Holes, leading to the Active Galactic Nuclei (AGN) phenomenon and relative feedback on the host galaxy. We compare gaea predictions based on the PMS , with model realizations based on other simulations in the Millennium Suite at different resolution, showing that the new model provides a remarkable consistency for most statistical properties of galaxy populations. We interpret this as due to the interplay between AGN feedback and H2-based SFR (both acting as regulators of the cold gas content of model galaxies), as model versions considering only one of the two mechanisms do not show the same level of consistency. We then compare model predictions with available data for the galaxy 2-point correlation function (2pCF) in the redshift range 0¡z3. We show that gaea runs correctly recover the main dependencies of the 2pCF as a function of stellar mass (M⋆), star formation activity, HI-content and redshift for M⋆ ¿ 109 M⊙ galaxies. These results suggest that our model correctly captures both the distribution of galaxy populations in the Large Scale Structure and the interplay between the main physical processes regulating their baryonic content, both for central and satellite galaxies. The model predicts a small redshift evolution of the clustering amplitude, that results in an overprediction of z3 clustering strength with respect to the available estimates, but is still consistent with data within 1- uncertainties.
@@
According to the classical interpretation of the 2pCF shape, it can be seen as a combination of (see, e.g., Cooray & Sheth, 2002): a 2-halo term - corresponding to pairs of galaxies belonging to independent halos - and a 1-halo term - coming from pairs of galaxies living in the same halo. The 2-halo term is dominant at large scales (separations larger than 3 Mpc) and comes mainly from pairs of central galaxies. The 1-halo term is responsible for the 2pCF shape at 2 Mpc and it is statistically dominated by galaxy pairs belonging to the same halo (that is at least one of the two galaxies is a satellite). Observationally, only the projected 2pCF - - is directly accessible, while the tridimensional 2pCF can be recovered via de-projection. Early results by Davis & Geller (1976) already showed a clear dependence of clustering amplitude with galaxy type, by showing that early type galaxies are more clustered than late-type galaxies. The advent of large area redshift surveys like the Sloan Digital Sky Survey (SDSS) provided the statistics to explore the dependence of amplitude and shape on galaxy properties such as luminosity (see, e.g., Zehavi et al., 2002) or stellar mass (Li et al., 2006), showing that brighter/more massive galaxies are more clustered than their fainter/less massive counterparts. A stronger difference in the clustering signal is found when considering galaxies of different colour (see, e.g., Li et al., 2006); at fixed M⋆ (or luminosity), red galaxies sensibly more clustered than blue galaxies. Galaxy colour is sensitive to the star formation rate (SFR) of model galaxies, which is regulated by a complex network of internal and external physical mechanisms. Therefore, clustering predictions for the active and passive galaxy populations represent a additional diagnostic to assess whether a given model reproduces the observational data in terms of star formation quenching as a function of the environment.
@@
Most of our previous gaea realizations have been run on dark matter merger trees extracted from the Millennium Simulation (Springel et al., 2005, MS hereafter). This realization is based on a cosmological model whose parameters are offset with respect to more recent measurements – particularly for the normalisation of the density fluctuations at the present day () and . In this paper, we will introduce a novel gaea realization, run on the planck millennium simulation (Baugh et al., 2019, PMS hereafter), featuring a bigger cosmological volume, an updated cosmology, a better mass resolution and a finer time sampling. In particular, we will discuss the redshift evolution of key statistical quantities like the GSMFs or the AGN LFs as well as of galaxy clustering, as a function of stellar mass and star formation activity.
@@ 2 N-body simulations and the galaxy formation model
0.073 | Star Formation efficiency | |
@@
rpstime | 444 | Timescale of hot gas ram-pressure stripping |
kesislow | 1.19 | Ratio between specific angular momentum of gas cooling through ‘slow mode’ and that of the halo |
kesirapid | 1.38 | Ratio between specific angular momentum of gas cooling through ‘rapid mode’ and that of the halo |
@@ 2.1 Numerical Simulations
In this work, we will focus on predictions from a gaea realization coupled with merger trees extracted from the PMS N-body simulation (Baugh et al., 2019). Moreover, we will compare our results against predictions from two other gaea runs, based on the millennium Simulation (Springel et al., 2005, MS) and millennium ii Simulation (Boylan-Kolchin et al., 2009, MSII). These runs corresponds to the predictions at the basis of our recent work (De Lucia et al., 2024; Xie et al., 2024). Details on the assumed cosmologies, box size and particle resolution for each N-body simulation are listed in Table 1.
@@
The initial conditions for the PMS have been generated at z=127 using second order Lagrangian perturbation theory (see Jenkins, 2013, for more details), and the simulation has been run using a reduced memory version of the N-body code gadget (Springel et al., 2005). Halo and Sub-Halo catalogues have been extracted on the fly using a classical FOF algorithm and the substructure finder code subfind (Springel et al., 2001): halos and sub-haloes information have been saved on a regular grid of 271 output times (i.e. snapshots), equally spaced in expansion factor. This has to be compared with the 64-68 snapshots of the MS and MSII over the same redshift range. The gain in temporal (redshift) resolution is particularly relevant at z4 where the PMS has 100 snapshots against the of the MS/MSII. In order to run GAEA on the PMS volume, we have constructed subhalo based merger trees using the same approach described in Springel et al. (2005). We refer to this paper for a detailed description of the algorithm. In summary, for any given subhalo, we identify a unique descendant by finding all subhalos at the subsequent output that contain its most bound particles. We then give higher weight to those particles that are more tightly bound to the subhalo under consideration, and select as descendant the subhalo containing a larger fraction of the most bound particles. In this way, we aim at tracing the fate of the inner regions of the substructure, that might survive for longer time after it has been accreted onto a larger structure. Following (Springel et al., 2005), we search for descendants in the two following outputs to deal with small subhalos that fluctuate below and above the assumed detection threshold (20 bound particles). Once a unique descendant has been identified for all subhalos in the simulated volume and for all available outputs, links to all progenitors are automatically established, and merger trees can be built. We adopt the same merger tree structure defined for the MS suite and shown in Fig. 11 of (Springel et al., 2005). The merger trees adopted in this work have been constructed using half of the available outputs. While reducing the number of redshifts available, this choice reduces the memory and computational requirements of our GAEA runs. In a forthcoming paper (Cantarella et al., in preparation), we will show that results are not significantly different from those obtained using merger trees constructed considering all available snapshots.
@@ 4.1 Redshift evolution of the 2-point correlation function
We then consider the predicted correlation functions for our model universes at different redshifts. In the following, we will focus on the projected 2-point correlation function ; this choice makes the comparison with observational data straightforward, and does not have to rely on de-projection algorithms. We study the evolution of the projected correlation function using estimates from several surveys: going from low-z to higher redshifts we compare against results from SDSS (Li et al., 2006, z0.1, maximum integration scale222The maximum integration scale represents the finite upper limit used for the integration of the real-space correlation function. =40 ), GAMA (Farrow et al., 2015, z¡0.2, =40 ), VIPERS (Marulli et al., 2013, , =30 ), DEEP2 (Li et al., 2006, z0.9, =40 ), VVDS (Meneux et al., 2008, z0.9, =20 ), zCOSMOS (zCOSMOS Meneux et al., 2009, z0.9, =20 ) and VUDS (Durkalec et al., 2018, z2.7, =20 ). We collect all these observational measurements in Fig. 3 and we contrast them with the prediction from gaea realizations based on the MS (black dashed lines), the MSII (dotted lines) or the PMS (red solid lines). Where not visible, the dotted lines (MSII) overlap with the dashed lines (MSI). We do not show predictions from the MSII for bins corresponding to M10 M⊙, as there are too few galaxies above this mass limit to provide a statistical estimate for the 2pCF. We compute from gaea using the python corrfunc package (Sinha & Garrison, 2020); model galaxy samples have been extracted from the simulated volumes considering all galaxies in the stellar mass range indicated in the legend, i.e. we do not attempt to replicate selection criteria specific of the different surveys considered. In each plot (as well as in Fig. 4 and 6), model estimates have been computed using the same as in the observed samples we compare with, but for the z0.9 panel. In this case we show model predictions corresponding to =40 (i.e. the value used in the DEEP2 survey). In our analysis we include an estimate for the typical uncertainty on M⋆ by convolving the intrinsic predictions of the model with a log-normal error distribution of amplitude 0.25 dex. We check that our main conclusions do not depend on this choice: as predictions based on intrinsic M⋆ provide a consistent picture. We also explore the impact of uncertainties in the theoretical determination by computing the jackknife errors using the CosmoBolognaLib package (Marulli et al., 2016). We show these uncertainties for the PMS as the orange areas. We also assess the impact of cosmic variance on the comparison between data and model predictions: we randomly extract from the PMS 100 boxes with the same volume as the observational survey. The yellow areas in Fig. 3 highlight the range covered by these sub-volumes. Overall, jackknife errors are relevant at large separations and for the most massive galaxies. Cosmic variance is relevant at the highest redshift considered and for small surveyed volumes, where it leads to errors that are slightly larger than those based on jackknife at intermediate scales.
Overall, the level of agreement between theoretical and observed correlation functions is satisfactory over the entire redshift range considered. Models run on the MS, MSII and PMS agree well with each other, but for a slight tendency of the MS and MSII runs to predict systematically stronger correlations than PMS . The effect is more pronounced at small stellar masses and close separations, and we ascribe it to the larger value of used in the MS and MSII. At z (Fig. 3, upper left panel), theoretical are very close to SDSS constraints for . Above and below this mass range, model predictions slightly underestimate and over-estimate the measured clustering signal, respectively.
The lower normalization of the 2pCF for galaxies more massive than 10 is connected with the overprediction of the high-mass end of the z0 GSMF (top left panel of Fig. 2). In this mass range, our model galaxies lie above the Mvir versus M⋆ relation that can be inferred from data using a classical subhalo abundance matching technique (Behroozi et al., 2013; Moster et al., 2013), i.e. gaea massive galaxies tend to live in DMHs less massive than what can be inferred using halo occupation distribution statistics.
For galaxies in the lowest mass bin considered, the overprediction of the clustering amplitude is mainly driven by the mass distribution of satellites and centrals in the smallest resolved DM haloes, an effect already discussed in Wang et al. (2013a). These authors showed that to reproduce at the same time the GSMF and 2pCF in the framework of an halo occupation distribution model, satellite galaxies should be less massive than centrals, on average, at fixed stellar mass. They also considered results from two independent semi-analytic models that struggled to reproduce these results - a problem that is shared also by our GAEA model. It is worth stressing that gaea lacks an explicit treatment for tidal stripping of stars from satellite galaxies, a process that could move model predictions in the right direction and whose impact we plan to study in future work. Moreover, at scales 2 Mpc orphan galaxies (i.e. satellite galaxies whose substructure has fallen below the resolution limit of the simulation) represent a relevant population and their orbital evolution and merging times estimates might be incorrect (see e.g. De Lucia et al. 2010 for a discussion).
Similar conclusions hold at intermediate redshifts () for VIPERS (Fig. 3, upper right panel). The situation is slightly more complex at z0.9, where different surveys suggest slightly different normalization for . This could be due to the different surveyed volume or to the presence of overdensities (like in the case of the zCOSMOS sample, (Meneux et al., 2009)). Nonetheless, gaea predictions lie within the clustering measurements obtained from different surveys at z0.9 (Fig. 3, lower left panel) and the agreement with data is acceptable to the smallest separations measured. The predicted evolution of the normalization at z1 is rather small: in most cases can be well approximated by a single power law over a wide separation scale; gaea also captures the observed flattening of the relation at low-z for the smallest separations probed by the SDSS.
Moving to higher redshifts (z), a larger discrepancy of the ) normalization can be appreciated with respect to the VUDS data (Fig. 3, lower right panel): gaea predictions are still consistent with observational measurements within the uncertainties, but tend to lie systematically above the data. Cosmic variance can also have a non-negligible impact here (the VUDS volume correspond to 1.75 107 Mpc3 roughly 1/175th of the PMS ). Overall, we can conclude that gaea is able to satisfactory reproduce the observed evolutionary trends of the correlation strength up to z3.5. This represents a non-trivial success for our latest gaea model release.
@@ 4.2 Colour/SFR dependence of the 2-point correlation function
As we already discussed, several authors (see e.g. Wang et al. 2013b) pointed out that the clustering strength depends not only on , but also on the star formation activity of galaxy population, with red and quiescent galaxies being more clustered than their blue and star forming counterparts with similar . In Fig. 4, we compare model predictions with SDSS data at z0.1 (Li et al., 2006, =40 ) and data up to z1.0 (Coil et al., 2017, =40 ) obtained combining the PRIsm MUlti-object Survey (PRIMUS Coil et al., 2011) with DEEP2. At all redshifts, we consider model predictions corresponding to quiescent and star-forming subsamples, using separation thresholds based on their specific SFR (sSFR = SFR/M⋆), for consistency with our previous work (DL24).
At z0.1, model galaxies are classified using a sSFR= 10-11 yrs-1 separation. This is formally different than the choice for SDSS galaxies, which have been split according to their g-r colour: however, we checked that using a classifications based on galaxy colours (e.g. as in Springel et al. 2018) does not qualitatively change our main results. Overall, gaea realizations show a stronger clustering signal of red galaxies with respect to blue ones, in qualitative agreement with observations. In more detail, our models reproduce well the clustering strength for red galaxies over the entire mass range considered. This is particularly true for the PMS realization, while some residual overestimate of the clustering signal holds at the lowest masses, and smallest scales. It is worth stressing that this is the range of scales and stellar masses where our model samples are dominated by orphan galaxies (that are typically associated with low levels of SFR and red colors). The clustering strength of blue galaxies is also reasonably well reproduced for intermediate masses (i.e. in the 109.5 ¡ M⋆/[] ¡ 1011 range). Outside this range, the clustering of star-forming galaxies shows the same trends discussed for the global population, thus suggesting that this galaxy population is driving the tensions between predictions and data reported in Sec. 4.1 and Fig. 3. At higher redshift, we contrast clustering estimates from the PRIMUS and DEEP2 surveys with gaea predictions by splitting model galaxies into a quiescent and star forming subsamples, using the same (redshift-dependent) sSFR threshold employed in the data (Coil et al., 2017):
The agreement between data and model predictions in the single mass bin available (Log(M]) ) is very good, both for the red and blue populations.
The comparison presented in Fig. 4 indicates that gaea realizations reproduce qualitatively the relative clustering amplitude between the red and blue populations. In order to better highlight this we consider in Fig. 5 the ratio between the clustering amplitudes of the red and blue population at different redshifts for some of the available mass bins. Fig. 5 shows that GAEA predictions reproduce qualitatively the observed trends, with deviations due to the discrepancies discussed in relation to Fig. 4. The results presented in this section suggest that the treatment for the physical mechanisms responsible for quenching in gaea not only correctly reproduces the overall assembly and evolution of the quenched galaxy population (DL24, Xie et al. 2024), but it is also broadly consistent with their expected distribution in the Large Scale Structure. Therefore, this analysis represents a further positive test suggesting that gaea correctly captures the relative role of internal and environmental processes in shaping the galaxy properties in different environments.
@@
In the previous sections, we study the 2pCF of galaxies selected on the basis of their stellar content. A complementary analysis involves the study of the cold gas content of galaxies, through the distribution of HI selected galaxies. Gas distribution in galaxies at low redshift can be mapped using the 21 cm HI hyper-fine emission line. In particular, the Arecibo Fast Legacy ALFA Survey (ALFALFA; Giovanelli et al. 2005) samples more than 5000 galaxies out to z0.05, enabling the study of their clustering properties Papastergis et al. (2013), showed that HI selected galaxies exhibit a lower amplitude with respect to stellar mass selected samples. Guo et al. (2017) used the 70% complete ALFALFA catalogue and showed that the 2pCF of of HI-selected galaxies depends strongly on the HI mass, i.e. galaxies with higher HI reservoirs are more strongly clustered on scales above a few Mpc than their counterparts with lower HI content. We compare Guo et al. (2017) estimates (=20 ) at different HI-mass (MHI) thresholds with gaea predictions in Fig. 6; we find a general agreement in the clustering amplitude and, most importantly, in its relative evolution as a function of MHI. The larger discrepancies are seen at the high-MHI end, although data and predictions are consistent within the statistical errorbars. This result is similar to those based on M⋆ selected galaxies, although the reference samples refer to different model galaxy populations, as most massive galaxies at in gaea have low gas fractions.